OpenStudy (anonymous):

without graphing, determine whether the function represents exponential growth or exponential decay then find the y-intercept y=129(1.63)^x

OpenStudy (loser66):

growth

OpenStudy (loser66):

I am not with you@***[ISURU]***

OpenStudy (loser66):

it's growth just because the base of exponential is>1, I mean (1.63) >1 , the function is growth. If it is <1, then the function is decay. let do some calculation (1.63)^0 = 1 (1.63)^1 = 1.63 (1.63)^2 = 2.6569 as x up, the value of function up--> growth If the base is <1 , let say 0.2 (0.2)^0 = 1 (0.2)^1 = 0.2 (0.2)^2 = 0.04 and so on as x up, the value of function down--> decay

OpenStudy (anonymous):

what is the problem is as follows: f(x) = 2^-x ? would it just be a straight line? see, i'm horrible with math, so i'm probably wrong..i'm just curious.

OpenStudy (***[isuru]***):

srry for my mistake :( .. u r correct... thanks for clearing it out :)

OpenStudy (loser66):

everybody has mistake, me too. :)